OpenStudy (anonymous):

vertex of x^2 + 6x - 16 = 0

OpenStudy (anonymous):

The vertex of a parabola is right in the middle of the two x-intercepts. So solve first for those by factoring: x^2+6x-16 = (x+8)(x-2) = 0 So your intercepts occur at x = -8, and 2. Then the halfway point between -8 and 2 is -3. Therefore, the x-coordinate of the vertex is -3! Plug in -3 to find the y-coordinate and you're done :)

OpenStudy (anonymous):

wow what is wrong with \[-\frac{b}{2a}\]?